OpenStudy (cutiecomittee123):

The angle between vectors (2,3) and (-3,2) is____degrees How do I find the angle?

OpenStudy (loser66):

90

OpenStudy (cutiecomittee123):

How did you calculate this?

OpenStudy (cutiecomittee123):

@Loser66

OpenStudy (loser66):

slopes. Calculate slope of vector 1, namely m1 slope of vector 2, namely m2 you can see \(m_2 =-\dfrac{1}{m_1}\) That shows they are perpendicular.

OpenStudy (loser66):

Method2: if you want to use Linear algebra, just calculate \(\theta = arcos \dfrac{<v1, v2>}{||v1||||v2||}\)

OpenStudy (owen3):

the dot product is zero , therefore the vectors are perpindicular therefore the angle is 90

OpenStudy (cutiecomittee123):

Thanks for that! I am visual and youre description definitly helped me to put it into perspective:) @owen3